Sarah is 10 years older than her sister Emma. At one time Sarah was three
times as old as Emma. How old was Emma when Sarah was three times as old
times as old as Emma.

Answer by PuppyGirl76(1)

S = E + 10
Y = years before or after
S+Y = 3E+Y
E+10 + Y = 3E + Y
E = 5